archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    CLOSED Percent Complete Details The DSLAM in GOB75 is not responding Date Tuesday 05 May 2015 12 19PM Reason for closing Done Comments 2 Comment by OVH Tuesday 05 May 2015 12 18PM We ve opened a ticket with FT

    Original URL path: http://status.ovh.es/?do=details&id=9393 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    a result of task FS 13021 Sessions are currently operating on the second link We will continue with investigations on the network equipment Date Monday 13 April 2015 10 55AM Reason for closing Done Comments 1 Comment by OVH Monday 13 April 2015 10 53AM We isolated the LNS in Paris from the backbone between approximately 22h and 23h after 2 routers in Telehouse2 tasks 13021 and 13029 went down

    Original URL path: http://status.ovh.es/?do=details&id=9170 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details We re going to change the DNS IPs on the modems with Failover IPs during Monday Tuesday and Wednesday night No impact on customers expected Date Thursday 05 March 2015 10 57AM Reason for closing

    Original URL path: http://status.ovh.es/?do=details&id=8860 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Comments 6 Comment by OVH Thursday 26 February 2015 09 57AM We will start with can59 firmware uploaded we will reboot the DSLAM to upgrade the FPGA Comment by OVH Thursday 26 February 2015 09 58AM can59 Supervision cards are up Service cards are currently rebooting Comment by OVH Thursday 26 February 2015 09 59AM ron59 firmware uploaded We will reboot the DSLAM to upgrade the FPGA Comment by OVH

    Original URL path: http://status.ovh.es/?do=details&id=8837 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    CLOSED Percent Complete Details We will update our DNS servers This shouldn t affect customers Date Thursday 19 February 2015 16 28PM Reason for closing Done Comments 2 Comment by OVH Thursday 19 February 2015 16 28PM Update of dns2

    Original URL path: http://status.ovh.es/?do=details&id=8792 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details We noticed a power problem on DSLAM AMB33 and LOR33 The customers usi these DSLAMs are impacted We will investigate Date Thursday 19 February 2015 10 15AM Reason for closing Done Comments 1 Comment by

    Original URL path: http://status.ovh.es/?do=details&id=8790 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    05 February 2015 16 34PM Reason for closing Done Comments 3 Comment by OVH Wednesday 04 February 2015 18 24PM Out of the 5 connections that are down we ve remounted 3 from GlobalSwitch We just need to remount 2 to TH2 A flap on a connection was able to cause several uplinks downlinks on the pair 10 between 16 14 and 16 17 CEST We will keep the task

    Original URL path: http://status.ovh.es/?do=details&id=8678 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    from 0 00 we will update the ADSL equipment following Lille NRA LIB59 A 30 minute cut is expected after this intervention Date Tuesday 03 February 2015 01 40AM Reason for closing Done Comments 3 Comment by OVH Tuesday 03 February 2015 01 39AM Firmware uploaded we will reboot DSLAM to upgrade FPGA Comment by OVH Tuesday 03 February 2015 01 39AM Supervisory card is up card services are currently

    Original URL path: http://status.ovh.es/?do=details&id=8670 (2015-10-15)
    Open archived version from archive



  •